OpenStudy (laddiusmaximus):

If F(x) = f(g(x)) where f(-2)=8 f'2(-2)=4 f'(5)=3 g(5)=-2 and g'(5)=6 find F'(5) I know there is a chain rule in here somewhere, I am confused on how to start this.

OpenStudy (anonymous):

f(-2)=8, f'2(-2)=4 , f'(5)=3, g(5)=-2 and g'(5)=6 What do you mean by f'2(-2)=4?

OpenStudy (laddiusmaximus):

sorry typo. its f'(-2)=4

OpenStudy (anonymous):

F(x)= f(g(x)) F'(x)= f'(g(x)) * g'(x) F'(5)= f'(g(5)) * g'(5) = f'(-2) * 6 = 4*6 =24 . Is the answer correct? a bit rusty on this front